OpenStudy (anonymous):

A computer with a monitor costs $899. If the computer costs $241 more than the monitor, what is the price of the computer? what is the price of the monitor?

OpenStudy (anonymous):

So you could start by x + y = 899 meaning x is computer and y is monitor thus adding up to 899 Then it says the computer is $241 more than the monitor so we can say that x = y + 241 So if you substitute the y + 241 into x (from x + y = 899) you get y + 241 + y = 899 So now just solve for y or the monitor 2y = 658 y = 329 So the price of the monitor is $329 Then plug 329 back into the first equation x + 329 = 899 Find x x = 570 so the price of the computer is $570. The way we can check this is by subtracting the computer of $241 so get the price of the monitor which is $329. Hope this helps, and good luck :)